Problem 20.26 m Review Part A What is the value of q? Express your answer using two significant figures. The electric potential 2.0 m from a point charge g is 3.7x104 V Submit Request Answer Provide Feedback
Answer #1

formula for potential due to point charge is

V = kq/r

q= Vr/k

= 3.7* 10^4 V(2.0)/9* 10^9

=8.2* 10^-6 C

= 8.2 uC